40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13949     Entry time: Tue Jun 12 14:47:37 2018
Author: gautam 
Type: Bureaucracy 
Category: General 
Subject: Unlabelled components from EX moved to SP table and labelled 

Steve mentioned two unlabelled optics were found at EX, relics from the Endtable upgrade.

  • One was a 1" 45 deg p-pol optic (Y1-1025-C-45P), it looks a bit scratched.
  • The other was a Beam Sampler (BSF10-C).

These are now labelled and forked down on the SP table.

ELOG V3.1.3-